OpenStudy (anonymous):

a particle starts from rest and travels horizontally with a constant acceleration...at some point it hits a wall where its velocity and acceleration are reversed without any change in their magnitude..if the paticle returns to the initial point find the ratio of the time taken for the forward and reverse journey??

OpenStudy (unklerhaukus):

so the velocity[m/s] is the the same magnitude before and after collision,

OpenStudy (unklerhaukus):

the distance is the same, hence the time is also the same

OpenStudy (unklerhaukus):

the ratio must be one

OpenStudy (anonymous):

the answer is root2 +1

OpenStudy (unklerhaukus):

doesn't that imply that the particle has slowed down?

OpenStudy (anonymous):

i don't think that it implies that the particle has slowed down because the particle is accelerating, so the velocity which it will posses at the time of collision would be different from the initial velocity.

OpenStudy (unklerhaukus):

oh yeah is has a constant acceleration

OpenStudy (anonymous):

can someone solve fot it

OpenStudy (ajprincess):

for the forward motion say the time taken is t1 and say the distance is s. s=1/2at1^2 v=at1 in both the equations u=0 because it starts from rest. for the reverse motion say the time taken is t2. the distance will be equal. so s=vt2+1/2at2^2 =at1t2+1/2at2^2 1/2at1^2=(at1t2+1/2at2^2) t1^2=2t1t2+t2^2 t1^2-2t1t1-t2^2=0 t1^2-2t1t2+t2^2=2t2^2 (t1-t2)^2=2t2^2 t1-t2=sqrt2t2 t1=sqrt2t2+t2 t1=t2(sqrt2+1) t1/t2=sqrt2+1

OpenStudy (ajprincess):

is that k?